OpenStudy (anonymous):

8. There are 9 candidates running for 3 seats on a committee. How many different election results are possible? (1 point) • 84 • 504 • 56

OpenStudy (anonymous):

84

OpenStudy (anonymous):

Thank you I'm done this quiz finally

OpenStudy (aaronq):

in case you were wondering how the person came to the answer, its \(\huge(^n_r)=\dfrac{n!}{r!(n-r)!}=(^9_3)=\dfrac{9!}{3!(9-3)!}\)